OpenStudy (anonymous):

Solving quadratic inequalities 22x2 + 16x = 64 < 0

OpenStudy (anonymous):

I think your question should be like 22x2 + 16x - 64 < 0. Am I right? @pastorshall

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@pastorshall OK so wait

OpenStudy (anonymous):

+64 +64?

OpenStudy (anonymous):

haven't done these in a while I forgot

OpenStudy (anonymous):

\[22x^2 + 16x - 64 < 0 \] \[11x^2 + 8x -32 < 0 \](Dividing both sides by 2) comparing it with \[ax^2 + bx +c < 0\] we find a= 11, b= 8, c=-32 Now let us use quadratic formula to solve this quadratic inequality. \[ \frac{-b \pm \sqrt{b^2-4ac}}{2a}<0\] \[ \frac{-8 \pm \sqrt{8^2-4 \times 11 \times (-32)}}{2 \times 11}<0\] \[ \frac{-8 \pm \sqrt{64+1408}}{22}<0\] \[ \frac{-8 \pm \sqrt{1472}}{22}<0\] Now you can solve onward it

OpenStudy (anonymous):

I am looking back at the problem and it shows that it is 22x2 + 16x +64 < 0 But that just can't be right, could it??

OpenStudy (anonymous):

not - 64

OpenStudy (anonymous):

I wonder if that is a typo?